Eliminate the parameter to write y as a function of x:

x = √t

y = 2t - 3



1
Expert's answer
2021-03-30T16:24:03-0400

Solution.

"x=\sqrt{t} \\
 y=2t-3."

"t=x^2," then "y=2t-3=2x^2-3."

Answer. "y=2x^2-3."